OpenStudy (anonymous):

find x intercepts of the function 4x^2 +8x-32

OpenStudy (wolf1728):

The x-intercepts of the function are the solutions of the equation.

OpenStudy (anonymous):

4x^2 + 8x -32 = 0 divide through by 4 x^2 + 2x - 8 = 0 (x+4)(x-2) x=-4 and x=2 x={-4,2}

OpenStudy (wolf1728):

You could also use the quadratic formula for the solutions x = [-b +-sqrt (b² - 4ac] / 2a for this equation we would use a = 4 b = 8 and c = -32 or we could use a = 1 b = 2 and c = -8